the car accelerates uniformly from rest to speed of 15km/h in 10s. find
(a) the acceleration in miter per second squire
(b) the distance covered during this period in "m"
1
Expert's answer
2019-03-04T09:03:07-0500

The acceleration

a=∆v/∆t=(4.17 m/s)/(10 s)=0.42 m/s^2

The distance

S=v_i t+(at^2)/2=0×10+(0.42×〖10〗^2)/2=21 m